OpenStudy (anonymous):

Can someone please help me? Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x 1 + sec2x sin2x = sec2x [(sin(x))/(1-cos(x))]+[(sin(x))/(1+cos(x))]=2csc(x) - tan2x + sec2x = 1

OpenStudy (anonymous):

I know I should start with the left side on the first one because it will be easier to manipulate.

OpenStudy (anonymous):

I think (sin x)(tan x cos x- cot x cos x)= (SOMETHING)(sinx/cosxSOMETHING)

OpenStudy (cruffo):

Good idea to start on the left.

OpenStudy (cruffo):

(sin x)(tan x cos x - cot x cos x) distribute tanx cosx sinx - cotx cosx sinx recall tanx = sinx / cosx and cotx = cosx/sinx

OpenStudy (anonymous):

ok so then i have \[\frac{ \sin x }{ \cos x } \cos x \sin x - \frac{ \cos x }{ \sin x} \cos x \sin x\] thanks for replying btw

OpenStudy (cruffo):

reduce to get sin^2x - cos^2x

OpenStudy (anonymous):

\[\frac{ \sin x }{ \cos x} \sin ^{2}x - \frac{ \cos x }{ \sin x} \cos ^{2}x\]

OpenStudy (anonymous):

like that?

OpenStudy (cruffo):

not quite... \[ \large \frac{ \sin x }{ \color{red} {\cancel {\cos x} }} \color{red} {\cancel{\cos x}} \sin x - \frac{ \cos x }{ \color{red} {\cancel{\sin x}}} \cos x \color{red} {\cancel{\sin x}}\]

OpenStudy (anonymous):

oh, ok. so then we have \[\sin ^{2}x-\cos ^{2}x\]

OpenStudy (cruffo):

We are pretty close to the end... Just to verify befor going on... Is the RHS \[\large 1 - 2 \cos(2x)\] or is it \[\large 1 - 2\cos^2(x)\]

OpenStudy (cruffo):

It's probably the second one.... I'm going to log off soon so I'll post the rest. Use Double Angle Formulas \[\large \cos(2x) = \cos^2x - \sin^2x = 2\cos^2x - 1\] so \[\large \sin^2x - \cos^2x \quad \color{green}{\text{factor out} -1}\] \[\large = -1\left(\cos^2x - \sin^2x\right) \quad \color{green}{\text{use above identity}}\] \[\large = -1\left(2\cos^2x - 1\right) \quad \color{green}{\text{distribute}}\] \[\large = -2\cos^2x + 1 \quad \color{green}{\text{rewrite}}\] \[\large = 1 -2\cos^2x \quad \color{green}{\text{done}}\]

OpenStudy (anonymous):

OMG, I actually understand that. Thanks so much.

OpenStudy (anonymous):

Totally makes sense!! I have been struggling for hours! I love you random citizen.

OpenStudy (anonymous):

@cruffo